40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 16 23:12:55 2012, kiwamu, Update, IOO, MC unlocking frequently MCsteps.png
    Reply  Fri Feb 17 01:02:22 2012, kiwamu, Update, IOO, MC is back to normal MCconnections.png
Message ID: 6292     Entry time: Fri Feb 17 01:02:22 2012     In reply to: 6291
Author: kiwamu 
Type: Update 
Category: IOO 
Subject: MC is back to normal 

[Koji / Kiwamu]

 The MC is now back to normal. The beam pointing to the interferometer is good.

There were two different issues :

  • A mechanical mount was in the MC WFS path.
  • There were some loose connections in the SUS rack

 

Slid have we the position of the mechanical mount. Nicely the WFS beam go through now.

And also I pushed all the connectors associated with the MC SUS OSEMs in the SUS rack.

After pushing the connectors, the MC1 OSEM readouts dramatically changed, which actually more confused us.

As shown in the 3 hours trend below, the OSEM readouts have changed a lot (shown in the middle of the plot with arrows). Some bumps after the steps correspond to our alignment efforts.

MCconnections.png

Quote from #6291

The MC became crazy again.

 

ELOG V3.1.3-